OpenStudy (beccab003):

Identify the line of symmetry of the parabola defined by y = -2(x + 4)^2 + 6 x=6 x= -4 x= -6 x=4

Directrix (directrix):

Do you know the x-coordinate of the vertex of the parabola?

OpenStudy (beccab003):

It's -6?

OpenStudy (anonymous):

do you know the equation to find the axis of symmetry?

OpenStudy (beccab003):

No

OpenStudy (anonymous):

\[x=\frac{-b }{2a }\]

Directrix (directrix):

@BeccaB003 What value of x makes (x + 4) equal to 0?

OpenStudy (beccab003):

-4

Directrix (directrix):

Yes, so x = -4 is the equation of the axis of symmetry. Look at the attached graph.

Directrix (directrix):

@BeccaB003

OpenStudy (beccab003):

Yes, I'm here. I was looking at what you said. So I don't need to worry about the other parts of the equation as long as (x+4)=0?

Directrix (directrix):

First up, did you see the graph?

OpenStudy (beccab003):

yes

OpenStudy (anonymous):

@Directrix will you please help me after?

Directrix (directrix):

The graph gives it away. To answer your question, when the parabola's equation is written in this type form: y = -2(x + 4)^2 + 6, you can pick off the coordinates of the vertex. In this form x + 4 the x coordinate of the vertex is whatever makes x+4 = 0. For this task of finding the axis of symmetry, the other terms are not necessary when the equation is written in this form.

OpenStudy (beccab003):

okay. Thanks for helping me understand better. You are most helpful. :)

Directrix (directrix):

See attachment. That form is called the vertex form of a parabola. Read more here: http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php

OpenStudy (beccab003):

Thanks!

Directrix (directrix):

You are welcome.